2-(50) Find Is, Ic IVCE and the region of operation IM2 20012 2 200
Answer #1

3V 10 kn {0.2k IB+ Ic IB Ic B=200 + UBE ein Assume this is working Active region. VB <VE Cutt off segion VCE 70.28 Active reg3 2-3= 3-0.7= 0.2 IB + 0.2 8 IB + 10 IB IB (0.2 + 40 + 103) IB= 2.3 1040.2 IB = = 2.21 X10 2.21X163 6 А LB = 2.21 MA Icz ß TBThis is working ein Active region IB= 2.21 MA LC = 0.44 mA VCE = 2.910